OBSERVATION OF COMBINATION BANDS INVOLVING INTERMOLECULAR VIBRATIONS OF N_{2}O-N_{2}, N_{2}O-OCS AND N_{2}O-CO_{2} COMPLEXES USING AN EXTERNAL CAVITY QUANTUM CASCADE LASER

\underline{M. REZAEI}, S. SHEYBANI-DELOUI, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, Alberta T2N 1N4, Canada; A.R.W. McKELLAR, National Research Council of Canada, Ottawa, Ontario, K1A 0R6, Canada.

Spectra of the weakly-bound N_{2}O-CO_{2}, N_{2}O-OCS, and N_{2}O-N_{2} complexes in the region of the N_{2}O ν_{1} fundamental band (∼2224 cm^{−1}) are observed in a pulsed supersonic slit jet expansion probed with a quantum cascade laser. One new band is observed for each complex: two combination bands involving the intermolecular in-plane bending for N_{2}O-CO_{2} and N_{2}O-N_{2} complexes, and the out-of-plane torsional vibration for N_{2}O-OCS. The resulting intermolecular frequencies are 34.17,^{a} 17.11 and 22.33 cm^{−1} for N_{2}O-CO_{2}, N_{2}O-OCS, and N_{2}O-N_{2} complexes, respectively. The intermolecular vibrations provide clear spectroscopic data against which theory can be benchmarked. These results will be discussed, along with a brief introduction to our pulsed-jet supersonic apparatus which has been retrofitted by an infrared cw external-cavity quantum cascade laser (QCL) manufactured by Daylight Solutions. The QCL is used in the rapid-scan signal averaging mode. Although the repetition rate of the QCL is limited by its PZT scan rate, which is 100 Hz, we describe a simple technique to increase the effective repetition rate to 625 Hz. In addition, we have significantly reduced the long term frequency drift of the QCL by locking the laser frequency to the sides of a reference line.

aLimin Zheng, Soo-Ying Lee, Yunpeng Lu, and Minghui Yang, J. Chem. Phys. 138, 044302 (2013).